# Proactive Agents for Multi-Turn Text-to-Image Generation Under Uncertainty

### v1

@@ E.8 Entity Parser Prompt Instruction
20 2. If the prompt indicates an entity definitly exists in the image, assign a 1.0 probability.
@@
E.9 Attribute Parser Prompt Instruction
@@ 53. The output should be a json parse-able format:
E.10 Relation Parse Prompt Instruction
@@
E.11 Verbalization Prompt Instruction
@@
E.12 Merge Prompt Prompt Instruction
@@
E.13 $AICQ_{base}$ Prompt Instruction
@@
E.14 $AICQ_{B}$ Prompt Instruction
@@
E.15 HSA Question Prompt Instruction
@@
Appendix F Automated evaluation
@@
Appendix G Details on the agent interface
Below is a showcase of how users could interact with the belief graph and clarifications in a hypothesised interface, to better iterate their inputs, to reach higher a quality and satisfaction of outputs. This is a crudely hypothesised, intentionally simple interface for the sake of research, but could be iterated and improved upon in many ways depending on application and users.
1. Default state On load of the app, there would be a text prompt input and space for output images, as is common across typical T2I interfaces. There would also be space for the user to view either clarifications from the model, or a graph interface, as part of the overall “input” section as these would act as a further input for future model output iterations. See Figure 13 below as reference.
@@ Appendix G Details on the agent interface
2. Output images, with Clarifications Once the user has submitted the prompt and the model has responded, there would be a set of images, as initial outputs from the users prompt. Below the input prompt would be a set of “Clarifications” in its populated state. These clarifications would ask the user specific questions that would be necessary to increase the specificity of the prompt, for the model to get a more accurate results aligned to the users intention, or to help the user realise their intention. Options would be given of the highest probability options for each Clarification, but the user could also fill in a totally new option via a free text field. Once answered by selection or text input, the clarifications would be added to the above, primary prompt for regeneration when the user selects. See Figure 14 below as reference.
@@
3. Graph Entities & Attributes Instead of the clarifications, the user could select to instead view a Graph by clicking Tab above the clarifications themselves. This graph would be populated will all Entities from the prompt explicit and implicit visually defined differently (in this diagram by the dotted line surrounds implicit entities, but is a filled line when surrounding explicit entities). The graph layout will be structured, depicting relationships concentrically i.e. "on", "in" or "under" for example, will become child entities, and be displayed within the parent entities’ boundary. For example a ’Mug’ that has the relationship of ’on’ a ’Table’ entity, will sit within the boundary of ’Table’, as also would a ’Plate’ if that had the same child-parent relationship.
Below the Graph would also be a list of ’cards’ (i.e. boxed groups of information), one for each "explicit" or "implicit" entity. Within each card a user could see the status of implicit / explicit, and change this status to confirm or deny its presence. The user could also see a list of "attributes" associated to that entity, which the model has assumed. Each of these attributes could be changed by interacting with a list of alternatives via drop down. These lists are determined in terms of which items and order of items, based on the probability by which the model sees them, ordered with higest first. This probability would be made clear to the user to define the order by seeing the peercentage next to the label. See Figure 15 below as reference.
@@
4. Graph Relationships The user would also be able to change the state of the Graph and Cards, to instead focus on the relationships between entities, by toggling to "Relations". In this state the user would be able to focus on two specific entities (e.g. ’mug’ and ’table’), see the description of the relationship (e.g. ’the mug is sitting on the table’) and if desired change the relationship to an alternative (e.g. ’on’, changed to ’under’) via a drop down of options which the model determined as alternative options ordered by probability, as per attributes. See Figure Figure 16 below as reference.
@@
Appendix H Details on user studies for the agent interface
@@ H.0.1 Hypothesized Frustrations
    "Entities" are single items that are intended to be in the image e.g. "Cat" and "Ball", from "make a sketch of a Cat playing with a Ball"
@@
    "Prompt" means the text written to communicate the intended output image e.g. the sentence "make a sketch of a Cat playing with a Ball" is the "Prompt", also known as "Input"
@@
    "Iterations" are each set of different image outputs by the model, taken from a different input, or even the same input just regenerated
The question asked for each Frustration were: "Please score the below frustrations (or issues) that could be related to Text to Image AI Generation"."Rank in terms of how much they relate to your current usage, with your most commonly used model or app."
H.0.2 Hypothesized Features
@@
    Clarifications: The model would ask specific clarifying questions about uncertainties in the prompt. These details would then be incorporated into subsequent iterations. For example: "Is the cat playing with: 1. a ball of wool, or 2. a tennis ball?"
@@
    Graph of Prompt Entities: A visual representation of all entities in the prompt as a graph, allowing users to see and edit attributes of each entity. E.g., seeing that the model has assigned "round," "small," and "wooden" as attributes to "table" and allowing the user to change them to "square" and "metal."
@@
    Graph of Prompt Relationships: A visual representation of relationships between entities in the prompt, allowing users to see and edit these relationships. E.g., seeing that "donut" is "next to" "coffee" and allowing the user to change the relationship to "on top of."
@@
    "How likely this feature is to help your current workflow if you had it now?". With response options of: "Very unlikely to help", "Unlikely to help", "Could help", "Likely to help", "Very likely to help".
@@
    "How soon would this feature deliver value to your work?" with response options of: "Very soon / immediately", "Sometime, "Not very soon".
@@
H.0.3 Human Study Results

### v2

@@ D.5 Entity Parser Prompt Instruction
20 2. If the prompt indicates an entity definitely exists in the image, assign a 1.0 probability.
@@
D.6 Attribute Parser Prompt Instruction
@@ 53. The output should be a json parse-able format:
D.7 Relation Parse Prompt Instruction
@@
D.8 Verbalization Prompt Instruction
@@
D.9 Merge Prompt Prompt Instruction
@@
D.10 $AICQ_{base}$ Prompt Instruction
@@
D.11 $AICQ_{B}$ Prompt Instruction
@@
D.12 HSA Question Prompt Instruction
@@
Appendix E Automatic Evaluation
@@
E.1 DSG Evaluations
The DSG (Cho et al., 2024) metric is used to compare the ground truth caption and the generated caption. DSG is adapted to parse text prompts into Davidsonian scene graph using the released code. The plots in Figure 11 show the T2T DSG metrics per turn. The results in Figure 11 show that significant gains in performance come from using proactive multi-turn agents. The blue line on the graph represents the simplest baseline which directly uses a T2I model and performs no updates to the initial prompt p0. Note that all multi-turn agents far exceed the baseline T2I model on all benchmarks for the DSG metric.
@@ E.1 DSG Evaluations
E.2 Mitigation of T2I Errors
A limitation of the proposed pipeline is that its overall capability is constrained by the prompt-following abilities of the employed text-to-image (T2I) model. The proposed agent prototypes call off-the-shelf T2I models directly, treating T2I APIs as tools. This offers the benefit of seamlessly switching to better T2I models when they become available. In an effort to mitigate T2I errors caused by the model not following the prompt, we perform an experiment demonstrating that using agent-user QA pairs can improve T2I fidelity across a batch of N seeds.
The experiment design is as follows: Each QA pair from the agent-user dialogue is converted into a (yes/no) VQA question concerned with a single detail of the image. Then, using a VQA score metric with these new questions, we can remove erroneous images from a set of N seeds by filtering out images with low VQA scores. We perform this experiment on the DesignBench image-caption dataset.
The specific implementation is as follows:
  1. 1.
    Use the 30 ground truth (GT) prompts of DesignBench and generate 10 images from 10 different random seeds with Imagen.
  2. 2.
    Take average DINO (I2I) score for all images against GT image, this was found to be 0.7637.
  3. 3.
    Take the first 5 Q-A pairs from Ag2 turn each into a binary yes or no question for which the correct answer is yes.
  4. 4.
    Run the VQA scorer over all 10 images per caption.
  5. 5.
    Choose best image of the ten by selecting the image with highest score.
  6. 6.
    Take average DINO (I2I) score for best image against GT image: this was found to be 0.7838.
  7. 7.
    Calculate the $\delta$ between before and after filtering out images via agent QA pairs: $\delta$ = +.02
In conclusion, this experiment demonstrates that by using the QA pairs from the agent in combination with the VQA score, we can improve image fidelity by filtering out images that do not follow the prompt. T2I models do not always follow a prompt exactly. They can make small errors or ignore a single detail while retaining all others. This is an inherent bottleneck of our agents; however, our results show that using the QA pairs from the agent-user interaction history allows us to overcome this limitation.
@@ E.2 Mitigation of T2I Errors
Appendix F Details on the Agent Interface
Below is a showcase of how users could interact with the belief graph and clarifications in a hypothesised interface, to better iterate their inputs, to reach a higher quality and satisfaction of outputs. This is a crudely hypothesised, intentionally simple interface for the sake of research, but could be iterated and improved upon in many ways depending on application and users.
1. Default state
On load of the app, there would be a text prompt input and space for output images, as is common across typical T2I interfaces. There would also be space for the user to view either clarifications from the model, or a graph interface, as part of the overall “input” section as these would act as a further input for future model output iterations. See Figure 12 below as reference.
@@ 1. Default state
2. Output images, with Clarifications
Once the user has submitted the prompt and the model has responded, there would be a set of images, as initial outputs from the users prompt. Below the input prompt would be a set of “Clarifications” in its populated state. These clarifications would ask the user specific questions that would be necessary to increase the specificity of the prompt, for the model to get a more accurate results aligned to the users intention, or to help the user realise their intention. Options would be given of the highest probability options for each Clarification, but the user could also fill in a totally new option via a free text field. Once answered by selection or text input, the clarifications would be added to the above, primary prompt for regeneration when the user selects. See Figure 13 below as reference.
@@ 2. Output images, with Clarifications
3. Graph Entities & Attributes
Instead of the clarifications, the user could select to instead view a Graph by clicking Tab above the clarifications themselves. This graph would be populated with all Entities from the prompt explicit and implicit visually defined differently (in this diagram by the dotted line surrounds implicit entities, but is a filled line when surrounding explicit entities). The graph layout will be structured, depicting relationships concentrically i.e. "on", "in" or "under" for example, will become child entities, and be displayed within the parent entities’ boundary. For example a ’Mug’ that has the relationship of ’on’ a ’Table’ entity, will sit within the boundary of ’Table’, as also would a ’Plate’ if that had the same child-parent relationship.
Below the Graph would also be a list of ’cards’ (i.e. boxed groups of information), one for each "explicit" or "implicit" entity. Within each card a user could see the status of implicit / explicit, and change this status to confirm or deny its presence. The user could also see a list of "attributes" associated to that entity, which the model has assumed. Each of these attributes could be changed by interacting with a list of alternatives via drop down. These lists are determined in terms of which items and order of items, based on the probability by which the model sees them, ordered with highest first. This probability would be made clear to the user to define the order by seeing the percentage next to the label. See Figure 14 below as reference.
@@ 3. Graph Entities & Attributes
4. Graph Relationships
The user would also be able to change the state of the Graph and Cards, to instead focus on the relationships between entities, by toggling to "Relations". In this state the user would be able to focus on two specific entities (e.g. ’mug’ and ’table’), see the description of the relationship (e.g. ’the mug is sitting on the table’) and if desired change the relationship to an alternative (e.g. ’on’, changed to ’under’) via a drop down of options which the model determined as alternative options ordered by probability, as per attributes. See Figure Figure 15 below as reference.
@@ 4. Graph Relationships
Appendix G Details on User Studies for the Agent Interface
@@ G.1 Hypothesized Frustrations
    “Entities” are single items that are intended to be in the image e.g. “Cat” and “Ball”, from “make a sketch of a Cat playing with a Ball”
@@
    “Prompt” means the text written to communicate the intended output image e.g. the sentence “make a sketch of a Cat playing with a Ball” is the “Prompt”, also known as “Input”
@@
    “Iterations” are each set of different image outputs by the model, taken from a different input, or even the same input just regenerated
The question asked for each Frustration were: “Please score the below frustrations (or issues) that could be related to Text to Image AI Generation”.” Rank in terms of how much they relate to your current usage, with your most commonly used model or app.”
G.2 Hypothesized Features
@@
    Clarifications: The model would ask specific clarifying questions about uncertainties in the prompt. These details would then be incorporated into subsequent iterations. For example: “Is the cat playing with: 1. a ball of wool, or 2. a tennis ball?”
@@
    Graph of Prompt Entities: A visual representation of all entities in the prompt as a graph, allowing users to see and edit attributes of each entity. E.g., seeing that the model has assigned “round,” “small,” and “wooden” as attributes to “table” and allowing the user to change them to “square” and “metal.”
@@
    Graph of Prompt Relationships: A visual representation of relationships between entities in the prompt, allowing users to see and edit these relationships. E.g., seeing that “donut” is “next to” “coffee” and allowing the user to change the relationship to “on top of.”
@@
    “How likely this feature is to help your current workflow if you had it now?”. With response options of: “Very unlikely to help”, “Unlikely to help”, “Could help”, “Likely to help”, “Very likely to help”.
@@
    “How soon would this feature deliver value to your work?” with response options of: “Very soon / immediately”, “Sometime, “Not very soon”.
@@
G.3 Human Study Results
@@
Appendix H Details on Recruitment and Participant Consent
In this human user study participants were recruited with the following requirements:
  1. 1.
    Situated in North America and fluent in English.
  2. 2.
    An age of 18 or above.
  3. 3.
    Have experience (i.e. at least once in the past month) using Text to Image AI Generation tools (e.g. Mid Journey, Stable Diffusion, ImageFX, DALL·E, Adobe Firefly etc.), for any purpose (work, non-work, just for fun).
  4. 4.
    The survey must be completed on a desktop computer - not a mobile phone.
  5. 5.
    Participants were recruited in North America and were compensated based on industry standards.
An internal review of the human-study experiment was first created to determine potential risks and harm to participants and found it to be low risk and we were not collecting any personally identifiable information. The review found that there were two potential risks. The first was the risk of a study with experimental technology: T2I AI models are still under development and may produce inaccurate, biased, or potentially offensive content. Participants must use caution when relying on the model’s output and always exercise critical thinking. The second risk regarded potential biases. The model may reflect the biases present in the data it was trained on. These biases can manifest in various ways, such as favoring certain groups or stereotypes. By using this model, participants acknowledge that they have read and understand this disclaimer. They also agree to use the model responsibly and ethically.
In response to the findings of the internal review, risk mitigation was performed by serving generated images rather than having participants generate images themselves using the model. All image-prompt pairs and all generated images shown to users were pre-approved as a way to reduce risk and remove any overtly harmful or offensive content. Additionally a disclaimer and self consent form were placed at the beginning of the human-study survey to let participants know the type of questions they would be answering.